# Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inference

### v1

@@ 4.1.3 AGN feedback
Energy and momentum injection from accreting massive black holes provide a third physical channel capable of generating local departures from hydrodynamical equilibrium. Although AGN in dwarf galaxies have historically been difficult to identify – owing to weak emission signatures, uncertain bolometric corrections, and confusion with stellar processes (e.g. Reines2013; Baldassare2017) – both theory and observations now indicate that low-mass systems can indeed host active nuclei, and that even modest AGN episodes can imprint measurable disturbances on the surrounding interstellar medium (e.g. ManzanoKing2020; Koudmani2025) affecting galaxy evolution (e.g. Koudmani2022; ArjonaGalvez2024).
@@ 4.1.4 Gravitational asymmetries
The three mechanisms discussed so far – stellar feedback, self-gravitating clumps, and AGN activity – often operate within a broader structural context: large-scale gravitational asymmetries in the underlying potential. These asymmetries can arise from a variety of sources, including triaxial dark matter haloes (e.g. Hayashi2006), lopsided stellar mass distributions (e.g. Rix1995; Zaritsky2013), perturbations from orbiting subhaloes (Kazantzidis2008), massive gas clumps, or more violent events such as minor mergers (e.g. Peschken2020). In practice, gravitational asymmetries are difficult to disentangle from the other drivers because they can be both cause and consequence: a clump may deepen the potential in an off-centre fashion, generating streaming motions and torques; a minor merger can simultaneously deposit fresh gas, trigger a starburst, feed the AGN and distort the potential.
@@ 4.2 Accounting for the disequilibrium budget
Figure 6: Cumulative accounting of the disequilibrium budget in simulated dwarf galaxy discs. Top panel: Gas-mass-weighted distributions of $\varepsilon$ for the three physical drivers considered — AGN energy injection (yellow), stellar feedback (magenta), and self-gravitating clumps (purple) — plotted cumulatively on top of one another. The grey solid line shows the global distribution for the full sample. Overlaps between drivers are explicitly visualised: regions where AGN and stellar feedback coincide are hatched with ‘\\’ lines; overlaps between stellar feedback and self-gravitating clumps are hatched with ‘//’ lines. Colours in these overlaps follow a blending of the corresponding driver colours. For clarity, we omit the two remaining possible overlaps (AGN + clumps, and all three drivers simultaneously) as they contribute $\ll 1$ per cent of the total mass-weight. Bottom panel: Fractional contribution of each process to the global mass-weight in each $\varepsilon$ bin (same colour and hatch coding as above). This panel quantifies how each mechanism dominates, shares with, or yields to others across the disequilibrium spectrum. Vertical dashed lines in both panels mark the quasi-equilibrium range, $|\varepsilon|\leq 0.20$.
@@
In integrated terms, roughly $75$ per cent of the total gas mass in our sample lies outside the quasi-equilibrium range (i.e., $|\varepsilon|>0.2$). Of this ‘non-equilibrium’ component, the three drivers recover $\approx 85$ per cent for $|\varepsilon|>1$, but only $\approx 30$ per cent in the near-equilibrium regime ($|\varepsilon|\leq 0.2$) – a scale-dependent recovery fraction that matches the visual impression from Fig. 6. The remaining untagged mass, particularly at $0.2<|\varepsilon|<1$, most likely reflects transient processes and our omission in Fig. 6 of the asymmetric potential as a driver (Sec. 4.1.4), as well as mild turbulent, thermal, or gravitational fluctuations below our selection thresholds.
@@ 5 Classification of Disequilibrium Galaxies
Determining which galaxies yield rotation curves suitable for dynamical mass modelling remains a longstanding problem. Common observational quality cuts – symmetric large-scale velocity gradients, high ordered-to-random motion ratios ($V_{\rm rot}/\sigma$), or thin, axisymmetric disc morphologies – are widely employed (e.g. Trachternach2008; Oh2015; Pina2025), yet none guarantees hydrodynamical equilibrium, nor do they constitute a uniform, physically grounded standard across environments.
@@ 5.1 The centrifugal–hydrodynamical equilibrium link
Figure 7: Visual summary of dynamical state classification of simulated gaseous discs. Each column corresponds to one of the three $k$-means classes (Appendix B), colour-coded consistently across rows: Class 1 (green), Class 2 (blue), and Class 3 (red). Top row: Azimuthally averaged rotational velocity profiles, $\langle v_{\phi}\rangle_{\phi}(r)$, for all galaxies in each class. Middle row: Corresponding azimuthally averaged circular velocity profiles, $\langle v_{c}\rangle_{\phi}(r)$. Bottom row: Fractional error between the model-dependent reconstructed circular velocity, $v_{c,\rm model}(r,\phi)$, and the true circular velocity from the potential $v_{c}(r,\phi)$, shown as class-stacked medians with $50^{\mathrm{th}}$-percentile envelopes (both weighted) as a function of radius. Lines correspond to the centrifugal model (dark orange/dashed), the pressure-supported model (dark red/dash dotted), and the equilibrium model (dark green/dotted) – i.e., the same colour and line-style schemes as in Fig. 1. Despite similar $\langle v_{c}\rangle_{\phi}$ shapes across classes, the residuals with respect to $\langle v_{\phi}\rangle_{\phi}$ reveal distinct dynamical states, which are then reflected in the structure of the fractional error distributions across different models, including the hydrodynamical equilibrium-based, reinforcing the link between centrifugal balance and hydrodynamical equilibrium.
@@
We quantify this connection by characterising each galaxy with three complementary residual-based metrics computed from $\langle v_{\phi}\rangle_{\phi}$ and $\langle v_{c}\rangle_{\phi}$: a global integrated deviation ($L_{1}$ norm), a fractional pointwise deviation in amplitude (NRMSE-Amplitude), and a fractional pointwise deviation in slope (NRMSE-Slope). Applying $k$-means clustering in this three-dimensional metric space yields three statistically well-separated groups, interpreted as dynamical classes. For clarity, Appendix B presents two case-study galaxies – a stable and a disturbed disc – showing the resulting metrics, together with a visualisation of the 3D parameter space used by our $k$-means clustering algorithm to identify the three dynamical classes and further details.
@@ 5.1.1 Class 1 – Coherent Rotators, $29$ per cent of the sample (35/122)
These galaxies exhibit near-perfect alignment between $\langle v_{\phi}\rangle_{\phi}$ and $\langle v_{c}\rangle_{\phi}$, with median residuals $\approx 0$ per cent across most radii and only mild ($\approx-5$ per cent) under-rotation beyond $r\gtrsim 12$kpc. This bias largely vanishes when pressure or advective terms are included. Inside $\approx 4~\mathrm{kpc}$ – and especially within $\approx 2~\mathrm{kpc}$ – all models show significantly broadened error distributions, with $50^{\mathrm{th}}$-percentile ranges extending beyond the $\pm 10$ per cent threshold, which might be critical to discuss local features (or bumps) in the rotation curve and determine whether a halo is cuspy or cored (e.g. Sands2024). Even for dynamically coherent discs, therefore, the innermost kiloparsec remains a somewhat challenging regime for reliable mass modelling.
@@ 5.2 Galaxy-wide correlations with dynamical class
A clear trend is seen in total Hi mass, i.e. $M_{\mathrm{H\textsc{i}}}=M_{\mathrm{H\textsc{i}}}(r\leq R_{\rm ext,H\textsc{i}})$. Coherent rotators have the highest median $M_{\rm H\textsc{i}}$; this is consistent with the physical picture of more inertia yielding enhanced resistance to perturbative forces, and the maintenance of well-ordered discs. Perturbed rotators occupy an intermediate regime with broader scatter, indicating that moderately gas-rich systems can retain some rotational support despite disturbances. Slow rotators show systematically lower $M_{\rm H\textsc{i}}$, suggesting that reduced Hi content is strongly associated with disrupted kinematics. This correlation is evident in our narrow halo-mass-selected sample ($M_{200\mathrm{c}}$) but may be obscured in $V_{\rm max,~rot}$–selected samples, where lower rotation speeds in disturbed systems bias them out of the selection – i.e., the observationally accessible $V_{\rm max,~rot}\equiv\max(\langle v_{\phi}\rangle_{\phi})$ will be a biased estimate of $V_{\rm max}\equiv\max(\langle v_{c}\rangle_{\phi})$ for this specific class. Similar trends are seen in the NIHAO simulations (Dutton2019) and in the FIRE/FIRE-3 simulations (ElBadry2018; Sands2024), which find that lower-mass, Hi-poor galaxies are less rotationally supported and more strongly affected by non-circular motions. As noted in Dutton2019, this is confirmed observationally from the skewness of linewidths in bins of Hi mass in both ALFALFA (Haynes2018) and HIPASS (Koribalski2004) observations.
@@
We quantify the gas disc thickness with the minor-to-major axis ratio ($c/a$). Thin discs are notoriously difficult to capture accurately in cosmological hydrodynamical simulations (e.g. Agertz2011; Tacchella2019; Ludlow2023; Benavides2025). COLIBRE seems to fare better in this regard than past efforts: our sample includes $39/122$ galaxies with stellar discs thinner than $c/a=0.3$, and $7$ galaxies with $c/a\leq 0.2$. This still falls short of the frequency of (very) thin stellar discs in dwarf galaxies inferred from observed shape distributions, but is a substantial improvement over other models including TNG50, FIREbox and ROMULUS25 (Benavides et al. in preparation). This is encouraging: thin gas discs are a prerequisite to the formation of thin stellar discs, and the gas $c/a$ distributions for our three dynamical classes show a monotonic progression that parallels the trend with total Hi mass, indicating a tight connection between gas disc morphology and dynamical state. In vertical hydrostatic equilibrium, greater turbulence implies a thicker disc, and observationally $V_{\rm rot}/\sigma$ correlates tightly with disc aspect ratio (see sec. 5.3.1 of Pina2025). Here, coherent rotators are thin (median $c/a\sim 0.05$) with minimal scatter; perturbed rotators are thicker (median $\sim 0.15$) with a broader range; slow rotators are systematically thick (median $\sim 0.3$, none below $\sim 0.15$). These results suggest that $c/a$ is a strong morphological indicator of disequilibrium in this halo-mass regime, and that its diagnostic potential in rotation curve reliability assessments should be considered of high fidelity.
For stellar and dark matter shapes, the distributions overlap between classes but show a bias toward greater prolateness (lower intermediate-to-major axis ratio $b/a$) in perturbed and slow rotators. This trend is consistent with findings from DowningOman2023, who identified similar correlations between halo shape and dynamical perturbations in high-resolution APOSTLE dwarf galaxies. Theoretically, this connection is well motivated: a galaxy embedded in a prolate halo often resides in a non-axisymmetric potential, which naturally excites strong non-circular motions in the gas kinematics (Hayashi2006; Hayashi2007; Marasco2018; Oman2019) – a behaviour our classification is explicitly sensitive to. Moreover, when the gas disc is significantly misaligned with the plane defined by the halo’s intermediate and major axes, no stable configuration can be sustained (see, e.g., the supplementary videos in DowningOman2023), further reinforcing the link between halo shape and kinematic disequilibrium.
@@
These trends are both intuitive and well-aligned with existing literature. For instance, Pontzen2012 demonstrated, albeit under different subgrid prescriptions, that repeated momentum injection from supernova explosions can cause localised gas parcels to migrate substantially within the gravitational potential. This process induces rapid orbital fluctuations and delays the re-establishment of equilibrium in the disc; moreover, it aligns nicely with what we have shown and discussed in Sec. 4.1.1 and Sec. 4.2.
Furthermore, while AGN feedback has often been regarded as energetically unimportant in low-mass galaxies, recent work suggests that even modest AGN episodes can induce significant dynamical disturbances – see, e.g., the analysis in Koudmani2025 and the detailed discussion in their sec. 4.2. High-resolution simulations show that intermittent AGN outbursts can drive cycles of expansion and re-accretion in the central gas, generating potential fluctuations and delaying the re-establishment of equilibrium (e.g. Martizzi2013; Peirani2017) – i.e. a mechanism similar to what described in Pontzen2012 regarding supernova explosions [^7] As noted by Martizzi2013, this mechanism is effective only when the AGN-driven cycle of gas expansion and subsequent re-accretion occurs on timescales comparable to the local dynamical time. In Sec. 4.1.3 we indeed find $|\partial_{t}v_{r}|\sim|g_{r}|$ (i.e. $|\varepsilon|\sim 1$) in regions adjacent to an active AGN, indicating that the relevant gas motions occur on precisely the timescales required for this mechanism to operate efficiently.. Observational evidence reinforces this picture: ManzanoKing2020 reported a strong association between AGN activity and disturbed ionized-gas kinematics in dwarf galaxies, suggesting that AGN may play a more substantial role in shaping disc structure and rotation-curve morphology than previously assumed. This interpretation accords with our results, where slow rotators exhibit the highest AGN feedback signatures – with median energy proxies more than $1~\mathrm{dex}$ above those of perturbed and coherent rotators – as well as with the broader trends discussed in Sec. 4.1.3 and Sec. 4.2.

### v2

@@ 4.1.3 AGN feedback
Energy and momentum injection from accreting massive black holes provide a third physical channel capable of generating local departures from hydrodynamical equilibrium. Although AGN in dwarf galaxies have historically been difficult to identify – owing to weak emission signatures, uncertain bolometric corrections, and confusion with stellar processes (e.g. Reines et al., 2013; Baldassare et al., 2017) – both theory and observations now indicate that low-mass systems can indeed host active nuclei, and that even modest AGN episodes can imprint measurable disturbances on the surrounding interstellar medium (e.g. Manzano-King and Canalizo, 2020; Koudmani et al., 2025) affecting galaxy evolution (e.g. Koudmani et al., 2022; Arjona-Gálvez et al., 2024).
@@ 4.1.4 Gravitational asymmetries
The three mechanisms discussed so far – stellar feedback, self-gravitating clumps, and AGN activity – often operate within a broader structural context: large-scale gravitational asymmetries in the underlying potential. These asymmetries can arise from a variety of sources, including triaxial dark matter haloes (e.g. Hayashi and Navarro, 2006), lopsided stellar mass distributions (e.g. Rix and Zaritsky, 1995; Zaritsky et al., 2013), perturbations from orbiting subhaloes (Kazantzidis et al., 2008), massive gas clumps, or more violent events such as minor mergers (e.g. Peschken et al., 2020). In practice, gravitational asymmetries are difficult to disentangle from the other drivers because they can be both cause and consequence: a clump may deepen the potential in an off-centre fashion, generating streaming motions and torques; a minor merger can simultaneously deposit fresh gas, trigger a starburst, feed the AGN and distort the potential.
@@ 4.2 Accounting for the disequilibrium budget
Figure 6: Cumulative accounting of the disequilibrium budget in simulated dwarf galaxy discs. Top panel: Gas-mass-weighted distributions of $\varepsilon$ for the three physical drivers considered — AGN energy injection (yellow), stellar feedback (magenta), and self-gravitating clumps (purple) — plotted cumulatively on top of one another. The grey solid line shows the global distribution for the full sample. Overlaps between drivers are explicitly visualised: regions where AGN and stellar feedback coincide are hatched with ‘\\’ lines; overlaps between stellar feedback and self-gravitating clumps are hatched with ‘//’ lines. Colours in these overlaps follow a blending of the corresponding driver colours. For clarity, we omit the two remaining possible overlaps (AGN + clumps, and all three drivers simultaneously) as they contribute $\ll 1$ per cent of the total mass-weight. Bottom panel: Fractional contribution of each process to the global mass-weight in each $\varepsilon$ bin (same colour and hatch coding as above). This panel quantifies how each mechanism dominates, shares with, or yields to others across the disequilibrium spectrum. Vertical dashed lines in both panels mark the quasi-equilibrium range, $|\varepsilon|\leq 0.20$. Inset: The Venn diagram provides a schematic representation of how gas parcels tagged by the different physical drivers are subsets of the global gas population, illustrating their partial overlap rather than implying distinct or self-normalised distributions. To recover the distribution associated with a given driver, one should consider only the coloured regions uniquely associated with that driver and its direct combinations, rather than the cumulative stacked profile.
@@
In integrated terms, roughly $75$ per cent of the total gas mass in our sample lies outside the quasi-equilibrium range (i.e., $|\varepsilon|>0.2$). Of this ‘non-equilibrium’ component, the three drivers recover $\approx 85$ per cent for $|\varepsilon|>1$, but only $\approx 30$ per cent in the near-equilibrium regime ($|\varepsilon|\leq 0.2$) – a scale-dependent recovery fraction that matches the visual impression from Fig. 6. The remaining untagged mass, particularly at $0.2<|\varepsilon|<1$, most likely reflects transient processes and our omission in Fig. 6 of the asymmetric potential as a driver (Sec. 4.1.4), as well as mild turbulent, thermal, or gravitational fluctuations below our selection thresholds. We quantify the sensitivity of these recovery fractions to the adopted tagging apertures and thresholds in Appendix B, where we show that the qualitative trends and relative driver contributions are robust to factor-of-two variations in the selection criteria.
@@ 5 Classification of Disequilibrium Galaxies
Determining which galaxies yield rotation curves suitable for dynamical mass modelling remains a longstanding problem. Common observational quality cuts – symmetric large-scale velocity gradients, high ordered-to-random motion ratios ($V_{\rm rot}/\sigma$), or thin, axisymmetric disc morphologies – are widely employed (e.g. Trachternach et al., 2008; Oh et al., 2015; Mancera Piña et al., 2025), yet none guarantees hydrodynamical equilibrium, nor do they constitute a uniform, physically grounded standard across environments.
@@ 5.1 The centrifugal–hydrodynamical equilibrium link
Figure 7: Visual summary of dynamical state classification of simulated gaseous discs. Each column corresponds to one of the three $k$-means classes (Appendix C), colour-coded consistently across rows: Class 1 (green), Class 2 (blue), and Class 3 (red). Top row: Azimuthally averaged rotational velocity profiles, $\langle v_{\phi}\rangle_{\phi}(r)$, for all galaxies in each class. Middle row: Corresponding azimuthally averaged circular velocity profiles, $\langle v_{c}\rangle_{\phi}(r)$. Bottom row: Fractional error between the model-dependent reconstructed circular velocity, $v_{c,\rm model}(r,\phi)$, and the true circular velocity from the potential $v_{c}(r,\phi)$, shown as class-stacked medians with $50^{\mathrm{th}}$-percentile envelopes (both weighted) as a function of radius. Lines correspond to the centrifugal model (dark orange/dashed), the pressure-supported model (dark red/dash dotted), and the equilibrium model (dark green/dotted) – i.e., the same colour and line-style schemes as in Fig. 1. Despite similar $\langle v_{c}\rangle_{\phi}$ shapes across classes, the residuals with respect to $\langle v_{\phi}\rangle_{\phi}$ reveal distinct dynamical states, which are then reflected in the structure of the fractional error distributions across different models, including the hydrodynamical equilibrium-based, reinforcing the link between centrifugal balance and hydrodynamical equilibrium.
@@
We quantify this connection by characterising each galaxy with three complementary residual-based metrics computed from $\langle v_{\phi}\rangle_{\phi}$ and $\langle v_{c}\rangle_{\phi}$: a global integrated deviation ($L_{1}$ norm), a fractional pointwise deviation in amplitude (NRMSE-Amplitude), and a fractional pointwise deviation in slope (NRMSE-Slope). Applying $k$-means clustering in this three-dimensional metric space yields three statistically well-separated groups, interpreted as dynamical classes. For clarity, Appendix C presents two case-study galaxies – a stable and a disturbed disc – showing the resulting metrics, together with a visualisation of the 3D parameter space used by our $k$-means clustering algorithm to identify the three dynamical classes and further details.
@@ 5.1.1 Class 1 – Coherent Rotators, $29$ per cent of the sample (35/122)
These galaxies exhibit near-perfect alignment between $\langle v_{\phi}\rangle_{\phi}$ and $\langle v_{c}\rangle_{\phi}$, with median residuals $\approx 0$ per cent across most radii and only mild ($\approx-5$ per cent) under-rotation beyond $r\gtrsim 12$kpc. This bias largely vanishes when pressure or advective terms are included. Inside $\approx 4~\mathrm{kpc}$ – and especially within $\approx 2~\mathrm{kpc}$ – all models show significantly broadened error distributions, with $50^{\mathrm{th}}$-percentile ranges extending beyond the $\pm 10$ per cent threshold, which might be critical to discuss local features (or bumps) in the rotation curve and determine whether a halo is cuspy or cored (e.g. Sands et al., 2024). Even for dynamically coherent discs, therefore, the innermost kiloparsec remains a somewhat challenging regime for reliable mass modelling.
@@ 5.2 Galaxy-wide correlations with dynamical class
A clear trend is seen in total Hi mass, i.e. $M_{\mathrm{H\textsc{i}}}=M_{\mathrm{H\textsc{i}}}(r\leq R_{\rm ext,H\textsc{i}})$. Coherent rotators have the highest median $M_{\rm H\textsc{i}}$; this is consistent with the physical picture of more inertia yielding enhanced resistance to perturbative forces, and the maintenance of well-ordered discs. Perturbed rotators occupy an intermediate regime with broader scatter, indicating that moderately gas-rich systems can retain some rotational support despite disturbances. Slow rotators show systematically lower $M_{\rm H\textsc{i}}$, suggesting that reduced Hi content is strongly associated with disrupted kinematics. This correlation is evident in our narrow halo-mass-selected sample ($M_{200\mathrm{c}}$) but may be obscured in $V_{\rm max,~rot}$–selected samples, where lower rotation speeds in disturbed systems bias them out of the selection – i.e., the observationally accessible $V_{\rm max,~rot}\equiv\max(\langle v_{\phi}\rangle_{\phi})$ will be a biased estimate of $V_{\rm max}\equiv\max(\langle v_{c}\rangle_{\phi})$ for this specific class. Similar trends are seen in the NIHAO simulations (Dutton et al., 2019) and in the FIRE/FIRE-3 simulations (El-Badry et al., 2018; Sands et al., 2024), which find that lower-mass, Hi-poor galaxies are less rotationally supported and more strongly affected by non-circular motions. As noted in Dutton et al. (2019), this is confirmed observationally from the skewness of linewidths in bins of Hi mass in both ALFALFA (Haynes et al., 2018) and HIPASS (Koribalski et al., 2004) observations.
@@
We quantify the gas disc thickness with the minor-to-major axis ratio ($c/a$). Thin discs are notoriously difficult to capture accurately in cosmological hydrodynamical simulations (e.g. Agertz et al., 2011; Tacchella et al., 2019; Ludlow et al., 2023; Benavides et al., 2025). COLIBRE seems to fare better in this regard than past efforts: our sample includes $39/122$ galaxies with stellar discs thinner than $c/a=0.3$, and $7$ galaxies with $c/a\leq 0.2$. This still falls short of the frequency of (very) thin stellar discs in dwarf galaxies inferred from observed shape distributions, but is a substantial improvement over other models including TNG50, FIREbox and ROMULUS25 (Benavides et al. in preparation). This is encouraging: thin gas discs are a prerequisite to the formation of thin stellar discs, and the gas $c/a$ distributions for our three dynamical classes show a monotonic progression that parallels the trend with total Hi mass, indicating a tight connection between gas disc morphology and dynamical state. In vertical hydrostatic equilibrium, greater turbulence implies a thicker disc, and observationally $V_{\rm rot}/\sigma$ correlates tightly with disc aspect ratio (see sec. 5.3.1 of Mancera Piña et al., 2025). Here, coherent rotators are thin (median $c/a\sim 0.05$) with minimal scatter; perturbed rotators are thicker (median $\sim 0.15$) with a broader range; slow rotators are systematically thick (median $\sim 0.3$, none below $\sim 0.15$). These results suggest that $c/a$ is a strong morphological indicator of disequilibrium in this halo-mass regime, and that its diagnostic potential in rotation curve reliability assessments should be considered of high fidelity.
For stellar and dark matter shapes, the distributions overlap between classes but show a bias toward greater prolateness (lower intermediate-to-major axis ratio $b/a$) in perturbed and slow rotators. This trend is consistent with findings from Downing and Oman (2023), who identified similar correlations between halo shape and dynamical perturbations in high-resolution APOSTLE dwarf galaxies. Theoretically, this connection is well motivated: a galaxy embedded in a prolate halo often resides in a non-axisymmetric potential, which naturally excites strong non-circular motions in the gas kinematics (Hayashi and Navarro, 2006; Hayashi et al., 2007; Marasco et al., 2018; Oman et al., 2019) – a behaviour our classification is explicitly sensitive to. Moreover, when the gas disc is significantly misaligned with the plane defined by the halo’s intermediate and major axes, no stable configuration can be sustained (see, e.g., the supplementary videos in Downing and Oman, 2023), further reinforcing the link between halo shape and kinematic disequilibrium.
@@
These trends are both intuitive and well-aligned with existing literature. For instance, Pontzen and Governato (2012) demonstrated, albeit under different subgrid prescriptions, that repeated momentum injection from supernova explosions can cause localised gas parcels to migrate substantially within the gravitational potential. This process induces rapid orbital fluctuations and delays the re-establishment of equilibrium in the disc; moreover, it aligns nicely with what we have shown and discussed in Sec. 4.1.1 and Sec. 4.2.
Furthermore, while AGN feedback has often been regarded as energetically unimportant in low-mass galaxies, recent work suggests that even modest AGN episodes can induce significant dynamical disturbances – see, e.g., the analysis in Koudmani et al. (2025) and the detailed discussion in their sec. 4.2. High-resolution simulations show that intermittent AGN outbursts can drive cycles of expansion and re-accretion in the central gas, generating potential fluctuations and delaying the re-establishment of equilibrium (e.g. Martizzi et al., 2013; Peirani et al., 2017) – i.e. a mechanism similar to what described in Pontzen and Governato (2012) regarding supernova explosions [^7] As noted by Martizzi et al. (2013), this mechanism is effective only when the AGN-driven cycle of gas expansion and subsequent re-accretion occurs on timescales comparable to the local dynamical time. In Sec. 4.1.3 we indeed find $|\partial_{t}v_{r}|\sim|g_{r}|$ (i.e. $|\varepsilon|\sim 1$) in regions adjacent to an active AGN, indicating that the relevant gas motions occur on precisely the timescales required for this mechanism to operate efficiently.. Observational evidence reinforces this picture: Manzano-King and Canalizo (2020) reported a strong association between AGN activity and disturbed ionized-gas kinematics in dwarf galaxies, suggesting that AGN may play a more substantial role in shaping disc structure and rotation-curve morphology than previously assumed. This interpretation accords with our results, where slow rotators exhibit the highest AGN feedback signatures – with median energy proxies more than $1~\mathrm{dex}$ above those of perturbed and coherent rotators – as well as with the broader trends discussed in Sec. 4.1.3 and Sec. 4.2.